college algebra - Nine times the first two consecutive odd integers equals seven times the second. What are the integers?

Answer by stanbon(57282)


Nine times the first two consecutive odd integers equals seven times the second. What are the integers?
--------------
1st: 2x-1
2nd: 2x+1
---
Equation:
9(2x-1) = 7(2x+1)
18x-9 = 14x+7
4x = 16
x = 4
----
1st: 2x-1 = 7
2nd: 2x+1 = 9
=====================
Cheers,
Stan H.